# Compositional Difference-in-Differences for categorical outcomes

### v2

@@ 4.1. Impact of early voting law on distributions of voters’ support across parties
This table shows that the two groups have almost similar racial compositions. Combined with their geographic proximity and political alignment over the study period, this makes them well-suited for causal analysis. The outcome of interest is the voter’s support across three categories, Democrats, Republicans, and Others. I consider that from 1992 elections to 2008. These data are publicly available and consistently reported across states. I start by plotting the log-counts to assess the plausibility of parallel growths.
@@
Overall, the policy increased turnout by around $4.4\%$, demonstrating that early voting can mobilize more voters. This finding is consistent with the existing literature on the effects of early voting. All three categories experienced a statistically significant increase in votes, although the magnitude of the effect varies considerably. The Democratic Party saw a modest 5.5% increase, while the Republican Party experienced a smaller rise of about 2.1%. In contrast, support for Other parties, which includes third-party and independent candidates, grew by nearly 37.5%, representing a much larger relative gain. Importantly, because “Other” parties started from a much smaller base, even a modest absolute increase translates into a large percentage change. Focusing on the two major parties, the Democratic Party benefited more from the policy than the Republican Party. This aligns with the results in Berry et al. (2025), which also found that such policies tend to favor Democrats over Republicans. However, I exercise caution in interpreting these results, as the states I selected were already Democratic-leaning during this period.
The next table presents the results for the compositional treatment effects (CTT).
@@
Party Estimate 95% CI Lower 95% CI Upper
Democrat 0.3136 0.3119 0.3151
Republican 0.2998 0.2982 0.3012
Other 0.3867 0.3838 0.3898
@@
Table 4 presents the estimated CTT for the 2008 election. Each entry represents the post-treatment probability of selecting a given party, under a counterfactual scenario where, before treatment, all parties were equally likely (i.e., each had a baseline probability of $1/3$). The results show a clear reallocation of voter preference. After the treatment: the probability of voting for “Other” rises to 38.7%, up from the baseline of 33.3%. The probability of voting Republican falls to 30.0%. The probability of voting Democratic also declines slightly, to 31.4%, but remains higher than that of Republicans. Thus, while both major parties lose relative appeal, Republicans lose more ground than Democrats, and third-party candidates gain the most in relative terms. This suggests the treatment not only increased overall turnout but also reshaped the internal composition of voter choice, pulling support more toward alternative options. This finding is interesting because most prior work focuses exclusively on the two major parties. My analysis reveals that policy interventions can meaningfully affect the electoral relevance of minor parties, a dimension often overlooked in the literature. It is worth emphasizing that this shift is purely relative: the overall ranking of parties by actual vote share in the treated states remains unchanged (Democrats first, Republicans second, Others third). Hence, the treatment does not overturn the existing political hierarchy, but it does narrow the gap between major and minor parties in terms of voter consideration.
4.2. Impact of RGGI on electricity generation composition
In this section, I study how the Regional Greenhouse Gas Initiative (RGGI) affected the mix of electricity generation in participating states. RGGI, launched in 2009, is a cooperative, market-based program among Connecticut, Delaware, Maine, Maryland, Massachusetts, New Hampshire, New York, Rhode Island, and Vermont. Its goal is to cap and reduce carbon emissions from the power sector. The program works by setting a regional cap on total carbon dioxide emissions from power plants and issuing a limited number of tradable allowances. Each allowance gives a power plant the right to emit one short ton of carbon dioxide. States distribute these allowances through quarterly auctions. The revenue from these auctions is reinvested in energy efficiency, renewable energy, and other programs that benefit consumers. Evaluating the impact of the Regional Greenhouse Gas Initiative (RGGI) on electricity generation is an interesting causal question because it reveals whether the policy effectively shifts energy production toward cleaner sources and reduces carbon emissions. I analyze U.S. electricity generation data, which is publicly available on the U.S. Energy Information Administration’s website (https://www.eia.gov/electricity/data/state/). This dataset provides annual electricity generation figures in megawatts by energy source for each state. I consider the following four categories of energy sources: gas (combining natural gas and other gases), coal and oil (combining coal and petroleum), nuclear (corresponding solely to the nuclear category), and renewables (combining hydroelectric, conventional, solar thermal and photovoltaic, geothermal, wind, and wood and wood-derived fuel). For each state and year, I observe the number of megawatts generated by different sources, and I aggregate these into two groups: the treated group and the control group. In constructing the control group, I exclude states that either joined RGGI, are strongly affected by leakage through electricity market integration, or implemented their own carbon pricing policies. In particular, Pennsylvania, West Virginia, Virginia, and Ohio are excluded due to their geographic proximity and electricity market linkages with RGGI states, which expose them to leakage and anticipatory effects. In fact, electricity markets in the U.S. Northeast and Mid-Atlantic are highly integrated (PJM, NYISO, ISO-NE), including these states in the control group risks contamination. California, Washington, and Oregon are also excluded, since those states adopted cap-and-trade or cap-and-invest programs that constitute carbon pricing mechanisms similar in nature to RGGI. Including them in the control pool would compromise the validity of the counterfactual by mixing treated and untreated units.
Figure 9. Map of the United States showing the nine RGGI states in purple and six excluded control states (due to market leakage or carbon pricing policies) in dark gray, and the control group consists of the gray states.
The program was organized into three-year compliance periods, with the first spanning January 1, 2009, to December 31, 2011. I focus my analysis on this initial period because the parallel growth assumption becomes progressively less plausible over longer time horizons. Here, I also assess the validity of the parallel growths assumption by plotting both the raw quantities and their logarithmic transformations over the study period (2000-2011). As shown in Figure 10, the log-quantity trajectories for the treated and control groups move approximately in parallel during the pre-treatment period, supporting the plausibility of the assumption.
Figure 10. log-quantity evolution 2000-2011 (treated vs control group)
I examine the impact of RGGI over the first three years of implementation (2009–2011), corresponding to the program’s initial compliance period. As shown in the dynamic estimates below, the policy induced a statistically significant and growing reduction in total electricity generation (see figure11).
Figure 11. GTT on total electricity production
These results show an 8.2% decline in total electricity generation in the first year, increasing to 8.9% in the second year, and reaching 13.6% by the end of the initial compliance period in 2011. The growing effect suggests that RGGI not only reduced generation in the short term but also triggered accelerating adjustments over time, potentially through coal plant retirements, fuel switching, or improvements in demand-side efficiency.
While total output fell, the composition of generation shifted, as I can see in table 5.
Table 5. Observed and Counterfactual Shares Across Three Years
Year Category Observed Counterfactual
Estimate 2.5% 97.5% Estimate 2.5% 97.5%
2009 Gas 0.3179 0.3178 0.3179 0.3146 0.3130 0.3161
coal & oil 0.1954 0.1953 0.1954 0.2123 0.2113 0.2132
Nuclear 0.3107 0.3106 0.3107 0.2818 0.2802 0.2833
Renewables 0.1761 0.1761 0.1762 0.1914 0.1896 0.1932
2010 Gas 0.3594 0.3593 0.3594 0.3215 0.3199 0.3231
coal & oil 0.1815 0.1815 0.1816 0.2127 0.2117 0.2137
Nuclear 0.2982 0.2982 0.2983 0.2716 0.2701 0.2731
Renewables 0.1609 0.1608 0.1609 0.1942 0.1924 0.1960
2011 Gas 0.3927 0.3927 0.3928 0.3272 0.3257 0.3289
coal & oil 0.1322 0.1322 0.1322 0.1965 0.1956 0.1974
Nuclear 0.2955 0.2954 0.2955 0.2583 0.2568 0.2597
Renewables 0.1796 0.1795 0.1796 0.2180 0.2161 0.2199
Over the first three years of RGGI (2009–2011), coal and oil generation fell sharply, dropping from a counterfactual share of about 21% to just 13.2% by 2011, a decline of over 8 percentage points, or nearly 40% relative to the counterfactual. Interestingly, this reduction was not primarily offset by renewables, as one might expect, but by natural gas. Gas generation surged from a counterfactual 32.7% to an observed 39.3%, a 6.6-point increase, reflecting rapid fuel switching to a lower-carbon, flexible source. Nuclear output also rose above its counterfactual, increasing by 3.7 percentage points, suggesting that RGGI enhanced the economic value of existing zero-carbon baseload capacity. Renewables, in contrast, underperformed: despite national trends favoring wind and solar, their share in RGGI states remained below the counterfactual in all three years, ending 2011 at 18.0% versus a counterfactual of 21.8%. These patterns indicate that the early RGGI-driven transition was led by natural gas rather than renewables.
Overall, RGGI effectively reduced coal and oil dependence, achieving substantial emissions reductions, as reflected in a 13.6% decline in total generation, but primarily through fuel switching to gas and increased nuclear output, rather than new renewable deployment. This highlights the role of existing infrastructure in shaping the composition of the clean energy transition. It also reveals the importance of complementary policies, such as renewable subsidies, to accelerate the shift toward truly clean energy sources. The results are consistent with several other studies (see Yan (2021)).
@@ 5. Conclusion
This paper introduces compositional difference-in-differences (CoDiD), a causal inference framework for analyzing how treatments and policies affect entire categorical outcome quantities and distributions. It is especially suited to settings with discrete, unordered outcomes, such as employment status (employed, unemployed, out of the labor force), voting choice, or health categories, where the policy-relevant question is not whether a single average has changed, but how the composition of all categories has been reconfigured. CoDiD addresses a dual objective: it quantifies both absolute changes in total counts and relative shifts in shares (i.e., the probability mass function). This allows researchers and policymakers to identify which categories expanded or contracted, by how much, and in what combination of scale and composition. The framework is built on a parallel growths assumption, a multiplicative analog of standard parallel trends, which posits that, in the absence of treatment, the ratios of category sizes evolve proportionally over time. This assumption has a natural interpretation in random utility models: it corresponds to stable relative utility differences across categories, so that estimated effects map directly to shifts in underlying preferences. CoDiD is designed for practical use and naturally extends to richer settings, including relaxations of the identifying assumption, staggered treatment adoption, and a synthetic DiD analog for richer data contexts. A key direction for future work is the incorporation of continuous covariates, enabling adjustment for confounders and the estimation of heterogeneous treatment effects across subpopulations.
@@ A.1. Proof of Theorem 1
$\log\cdot(q_{1,1}^{N})-\log\cdot(q_{1,0}^{N})=\log\cdot(q_{0,1}^{N})-\log\cdot(q_{0,0}^{N}).$
Add $\log\cdot(q_{1,0}^{N})$ to both sides:
$\log\cdot(q_{1,1}^{N})=\log\cdot(q_{1,0}^{N})+\log\cdot(q_{0,1}^{N})-\log\cdot(q_{0,0}^{N}).$
Since all vectors are strictly positive (by Assumption 1), the logarithm is defined component-wise, and I can apply the component-wise exponential function $\exp\cdot$ to both sides of (1). Because $\exp\cdot$ and $\log\cdot$ are inverses on $\mathbb{R}_{>0}^{p}$, I obtain:
$q_{1,1}^{N}=\exp\cdot\left(\log\cdot(q_{1,0}^{N})+\log\cdot(q_{0,1}^{N})-\log\cdot(q_{0,0}^{N})\right).$
By definition (implied by the structure of the model and assumption 1, the scalar $S_{1,1}^{N}$ is the sum of the components of the vector $q_{1,1}^{N}$. Let the components be indexed by $c_{k}$, for $k=1,\dots,p$. Then:
$S_{1,1}^{N}=\sum_{k=1}^{p}q_{1,1}^{N}(c_{k})=\sum_{k=1}^{p}\exp\left(\log q_{1,0}^{N}(c_{k})+\log q_{0,1}^{N}(c_{k})-\log q_{0,0}^{N}(c_{k})\right),$
@@
$\pi_{1,1}^{N}=\frac{q_{1,1}^{N}}{S_{1,1}^{N}},$
where the division is component-wise (i.e., normalizing the vector $q_{1,1}^{N}$ to sum to 1). Since both numerator and denominator are identified in Steps 1 and 2, $\pi_{1,1}^{N}$ is identified as stated.
@@ A.2. Proof of proposition 1
$U^{N}_{gt}(c_{k})=\mu^{N}_{gt}(c_{k})-\log\left(\sum_{k}e^{\mu^{N}_{gt}(c_{k})}\right)+\log\left(S^{N}_{gt}\right)+\varepsilon_{gt}(c_{k})$
@@
$\pi_{gt}(c_{k})=\Pr\Big(U^{N}_{gt}(c_{k})\geq U^{N}_{gt}(c_{j})\text{ for all }j\neq k\Big).$
Since $\varepsilon_{gt}(c_{k})\sim\text{i.i.d. Gumbel}$, the choice probability has the closed-form logit formula:
$\pi_{gt}(c_{k})=\frac{\exp\Big(\mu^{N}_{gt}(c_{k})+\log(S^{N}_{gt})-\log(\sum_{j}e^{\mu^{N}_{gt}(c_{j})})\Big)}{\sum_{j}\exp\Big(\mu^{N}_{gt}(c_{j})+\log(S^{N}_{gt})-\log(\sum_{l}e^{\mu^{N}_{gt}(c_{l})})\Big)}.$
@@
$\pi_{gt}(c_{k})=\frac{e^{\mu^{N}_{gt}(c_{k})}\cdot\frac{S^{N}_{gt}}{\sum_{l}e^{\mu^{N}_{gt}(c_{l})}}}{\sum_{j}e^{\mu^{N}_{gt}(c_{j})}\cdot\frac{S^{N}_{gt}}{\sum_{l}e^{\mu^{N}_{gt}(c_{l})}}}=\frac{e^{\mu^{N}_{gt}(c_{k})}}{\sum_{j}e^{\mu^{N}_{gt}(c_{j})}}.$
The $S^{N}_{gt}$ term cancels out in the probability, which shows that the relative shares depend only on the relative utilities, while $S^{N}_{gt}$ scales the absolute number of choices. The number of individuals choosing $c_{k}$ is:
$q_{gt}^{N}(c_{k})=\pi_{gt}(c_{k})\cdot S^{N}_{gt}=\frac{e^{\mu^{N}_{gt}(c_{k})}}{\sum_{j}e^{\mu^{N}_{gt}(c_{j})}}\cdot S^{N}_{gt}.$
@@
$\mathbb{E}[U_{g1}(c_{k})]-\mathbb{E}[U_{g0}(c_{k})]$ $=$ $\mu^{N}_{g1}(c_{k})-\log(\sum_{k}e^{\mu^{N}_{g1}(c_{k})})+\log\left(S^{N}_{g,1}\right)+\gamma$
$-\mu^{N}_{g0}(c_{k})-\log(\sum_{k}e^{\mu^{N}_{g0}(c_{k})})+\log\left(S^{N}_{g0}\right)-\gamma$
$=$ $\log q_{g1}^{N}(c_{k})-\log q_{g1}^{N}(c_{k})$
@@
$\mathbb{E}[U^{N}_{11}(c_{k})]-\mathbb{E}[U^{N}_{10}(c_{k})]=\mathbb{E}[U^{N}_{01}(c_{k})]-\mathbb{E}[U^{N}_{00}(c_{k})],\quad\forall c_{k}\in\bar{\mathcal{Y}}.$
@@ A.3. Proof of Proposition 2
$\log\left(\frac{\pi_{g,t}^{N}(c_{k})}{\pi_{g,t}^{N}(c_{p})}\right)=\log\left(\frac{q_{g,t}^{N}(c_{k})}{S_{g,t}^{N}}\times\frac{S_{g,t}^{N}}{q_{g,t}^{N}(c_{p})}\right)=\log(q_{g,t}^{N}(c_{k}))-\log(q_{g,t}^{N}(c_{p})).$
@@
$\log(q_{1,1}^{N}(c_{p}))-\log(q_{1,0}^{N}(c_{p}))=\log(q_{0,1}^{N}(c_{p}))-\log(q_{0,0}^{N}(c_{p})).$
Taking the difference between the two equalities gives:
$\big\{\log(q_{1,1}^{N}(c_{k}))-\log(q_{1,1}^{N}(c_{p}))\big\}-\big\{\log(q_{1,0}^{N}(c_{k}))-\log(q_{1,0}^{N}(c_{p}))\big\}=$
$\big\{\log(q_{0,1}^{N}(c_{k}))-\log(q_{0,1}^{N}(c_{p}))\big\}-\big\{\log(q_{0,0}^{N}(c_{k}))-\log(q_{0,0}^{N}(c_{p})).\big\}$
@@
$\log\left(\frac{\pi_{1,1}^{N}(c_{k})}{\pi_{1,1}^{N}(c_{p})}\right)-\log\left(\frac{\pi_{0,1}^{N}(c_{k})}{\pi_{0,1}^{N}(c_{p})}\right)=\log\left(\frac{\pi_{1,0}^{N}(c_{k})}{\pi_{1,0}^{N}(c_{p})}\right)-\log\left(\frac{\pi_{0,0}^{N}(c_{k})}{\pi_{0,0}^{N}(c_{p})}\right).$
From part 1, I also can derive that:
$\big\{\log(q_{1,1}^{N}(c_{k}))-\log(q_{1,1}^{N}(c_{l}))\big\}-\big\{\log(q_{1,0}^{N}(c_{k}))-\log(q_{1,0}^{N}(c_{l}))\big\}=$
$\big\{\log(q_{0,1}^{N}(c_{l}))-\log(q_{0,1}^{N}(c_{l}))\big\}-\big\{\log(q_{0,0}^{N}(c_{k}))-\log(q_{0,0}^{N}(c_{l})).\big\}$
@@
$\Big\{\mathbb{E}[U^{N}_{11}(c_{k})]-\mathbb{E}[U^{N}_{11}(c_{l})]\Big\}-\Big\{\mathbb{E}[U^{N}_{10}(c_{k})]-\mathbb{E}[U^{N}_{10}(c_{l})]\Big\}=$
$\Big\{\mathbb{E}[U^{N}_{01}(c_{k})]-\mathbb{E}[U^{N}_{01}(c_{l})]\Big\}-\Big\{\mathbb{E}[U^{N}_{00}(c_{k})]-\mathbb{E}[U^{N}_{00}(c_{l})]\Big\},\quad\forall\,c_{k},c_{l}\in\bar{\mathcal{Y}}.$ (14)

### v3

@@ 3. Application: Impact of early voting law on distributions of voters’ support across parties
This table shows that the two groups have almost similar racial compositions. Combined with their geographic proximity and political alignment over the study period, this makes them well-suited for causal analysis. The outcome of interest is the voter’s support across three categories: Democrats, Republicans, and Others. I consider that from the 1992 elections to 2008. These data are publicly available and consistently reported across states. I start by plotting the log-counts to assess the plausibility of parallel growths.
@@
Overall, the policy increased turnout by around $4.4\%$, demonstrating that early voting can mobilize more voters. This finding is consistent with the existing literature on the effects of early voting. All three categories experienced a statistically significant increase in votes, although the magnitude of the effect varies considerably. The Democratic Party saw a modest 5.5% increase, while the Republican Party experienced a smaller rise of about 2.1%. In contrast, support for Other parties, which includes third-party and independent candidates, grew by nearly 37.5%, representing a much larger relative gain. Importantly, because “Other” parties started from a much smaller base, even a modest absolute increase translates into a large percentage change. Focusing on the two major parties, the Democratic Party benefited more from the policy than the Republican Party. This aligns with the results in Berry et al. (2025), which also found that such policies tend to favor Democrats over Republicans. However, I exercise caution in interpreting these results, as the states I selected were already Democratic-leaning during this period. The next table 3 gives us what happened with the shares.
Table 3. Estimated categorical distributions
Party Estimate [CI] Counterfactual[CI] ATT [CI]
Democrat 0.5915 [0.5911, 0.5915] 0.5823 [0.5818, 0.5829] 0.0092 [0.0086, 0.0099]
Republican 0.3959 [0.3954, 0.3962] 0.4076 [0.4071. 0.4082] -0.0117 [-0.0124, -0.0111]
Other 0.0126 [0.0125, 0.0127] 0.0101[0.0100, 0.0102] 0.0025 [0.0024, 0.0027]
Table 3 reports the observed distribution of voter support in the 2008 elections and the estimated counterfactual distribution implied by Theorem 1. The results suggest that the early voting policy slightly increased support for the Democratic Party relative to what we would expect in the absence of treatment: the share of Democratic support is estimated at $0.5915$ but would have been $0.5823$ without treatment. However, Republican support would have been higher in the absence of treatment ($0.4076$ counterfactual vs. $0.3959$ observed), indicating a shift away from Republicans toward Democrats, suggesting that the main absolute redistribution of probability mass occurred between the two major parties. This is consistent with some of the results obtained in Berry et al. (2025). Interestingly, support for third parties also increased, suggesting that the treatment may have mobilized some voters outside the two major parties as well. Compared to the major parties, the outside option experienced the largest increase relative to its initial share. The next table (4) presents the results for the compositional treatment effects (CTT).
@@
Party Estimate [95% CI]
Democrat 0.3136 [0.3119, 0.3151]
Republican 0.2998 [0.2982, 0.3012]
Other 0.3867 [0.3838, 0.3898]
@@
Each entry represents the post-treatment probability of selecting a given party, under a counterfactual scenario where, before treatment, all parties were equally likely (i.e., each had a baseline probability of $1/3$). After the treatment: the probability of voting for “Other” rises to 38.7%, up from the baseline of 33.3%. The probability of voting Republican falls to 30.0%. The probability of voting Democratic also declines slightly, to 31.4%, but remains higher than that of Republicans. Thus, while both major parties lose relative appeal, Republicans lose more ground than Democrats, and third-party candidates gain the most in relative terms. This finding is interesting because most prior work focuses exclusively on the two major parties. My analysis reveals that policy interventions can meaningfully affect the electoral relevance of minor parties, a dimension often overlooked in the literature. It is worth emphasizing that this shift is purely relative: the overall ranking of parties by actual vote share in the treated states remains unchanged (Democrats first, Republicans second, Others third). Hence, the treatment does not overturn the existing political hierarchy, but it does narrow the gap between major and minor parties in terms of voter consideration.
@@ 4. Conclusion
This paper introduces Compositional Difference-in-Differences (CoDiD), a causal inference framework for analyzing how treatments and policies affect entire vectors of categorical quantities and their distributions. CoDiD is particularly suited to settings with discrete, unordered outcomes, such as employment status (employed, unemployed, out of the labor force), voting choices, or health categories, where the policy-relevant question is not just whether an average has changed, but how the composition of all categories and the total quantity have evolved. The framework simultaneously captures absolute changes in totals and relative shifts in shares, allowing researchers and policymakers to identify which categories expanded or contracted and by how much, in both scale and composition. Identification relies on a parallel growth assumption, a multiplicative analog of parallel trends, which posits that, absent treatment, category ratios evolve proportionally over time. This assumption has a natural behavioral interpretation in random utility models, mapping estimated effects to shifts in underlying preferences. CoDiD is practical and flexible, extending to settings with relaxed identifying assumptions. Future work may incorporate continuous covariates to adjust for confounding and estimate heterogeneous treatment effects across subpopulations as well as staggered treatment adoption.
@@ A.1. Proof of Theorem 1
$\log(q_{1,1}^{0})-\log(q_{1,0}^{0})=\log(q_{0,1}^{0})-\log(q_{0,0}^{0}).$
Add $\log(q_{1,0}^{0})$ to both sides:
$\log(q_{1,1}^{0})=\log(q_{1,0}^{0})+\log(q_{0,1}^{0})-\log(q_{0,0}^{0}).$
Since all vectors are strictly positive (by Assumption 1), the logarithm is defined component-wise, and I can apply the component-wise exponential function $\exp$ to both sides of (1). Because $\exp$ and $\log$ are inverses, I obtain:
$q_{1,1}^{0}=\exp\left(\log(q_{1,0}^{0})+\log(q_{0,1}^{0})-\log(q_{0,0}^{0})\right).$
By definition (implied by the structure of the model and assumption 1, the scalar $S_{1,1}^{0}$ is the sum of the components of the vector $q_{1,1}^{0}$. Let the components be indexed by $c_{k}$, for $k=1,\dots,p$. Then:
$S_{1,1}^{0}=\sum_{k=1}^{p}q_{k,1,1}^{0}=\sum_{k=1}^{p}\exp\left(\log q_{k,1,0}^{0}+\log q_{k,0,1}^{0}-\log q_{k,0,0}^{0}\right),$
@@
$\pi_{1,1}^{0}=\frac{q_{1,1}^{0}}{S_{1,1}^{0}},$
where the division is component-wise (i.e., normalizing the vector $q_{1,1}^{0}$ to sum to 1). Since both numerator and denominator are identified in Steps 1 and 2, $\pi_{1,1}^{0}$ is identified as stated.
@@ A.2. Proof of proposition 1
$U^{0}_{k,g,t}=\mu^{0}_{k,g,t}-\log\left(\sum_{k}e^{\mu^{0}_{k,g,t}}\right)+\log\left(S^{0}_{g,t}\right)+\varepsilon^{0}_{k,g,t}$
@@
$\pi_{k,g,t}=\Pr\Big(U^{0}_{k,g,t}\geq U^{0}_{j,g,t}\text{ for all }j\neq k\Big).$
Since $\varepsilon^{0}_{k,g,t}\sim\text{i.i.d. Gumbel}$, the choice probability has the closed-form logit formula:
$\pi_{k,g,t}=\frac{\exp\Big(\mu^{0}_{k,g,t}+\log(S^{0}_{g,t})-\log(\sum_{j}e^{\mu^{0}_{j,g,t}})\Big)}{\sum_{j}\exp\Big(\mu^{0}_{j,g,t}+\log(S^{0}_{g,t})-\log(\sum_{l}e^{\mu^{0}_{l,g,t}})\Big)}.$
@@
$\pi_{k,g,t}=\frac{e^{\mu^{0}_{k,g,t}}\cdot\frac{S^{0}_{g,t}}{\sum_{l}e^{\mu^{0}_{l,g,t}}}}{\sum_{j}e^{\mu^{0}_{j,g,t}}\cdot\frac{S^{0}_{g,t}}{\sum_{l}e^{\mu^{0}_{l,g,t}}}}=\frac{e^{\mu^{0}_{k,g,t}}}{\sum_{j}e^{\mu^{0}_{j,g,t}}}.$
The $S^{0}_{g,t}$ term cancels out in the probability, which shows that the relative shares depend only on the relative utilities, while $S^{0}_{g,t}$ scales the absolute number of choices. The number of individuals choosing $c_{k}$ is:
$q_{k,g,t}^{0}=\pi_{k,g,t}\cdot S^{0}_{g,t}=\frac{e^{\mu^{0}_{k,g,t}}}{\sum_{j}e^{\mu^{0}_{j,g,t}}}\cdot S^{0}_{g,t}.$
@@
$\mathbb{E}[U^{0}_{k,g,1}]-\mathbb{E}[U^{0}_{k,g,0}]$ $=$ $\mu^{0}_{k,g,1}-\log(\sum_{k}e^{\mu^{0}_{k,g,1}})+\log\left(S^{0}_{g,1}\right)+\gamma$
$-\mu^{0}_{k,g,0}-\log(\sum_{k}e^{\mu^{0}_{k,g,0}})+\log\left(S^{0}_{g,0}\right)-\gamma$
$=$ $\log q_{k,g,1}^{0}-\log q_{k,g,0}^{0}$
@@
$\mathbb{E}[U^{0}_{k,1,1}]-\mathbb{E}[U^{0}_{k,1,0}]=\mathbb{E}[U^{0}_{k,0,1}]-\mathbb{E}[U^{0}_{k,0,0}],\quad\forall k$
@@ A.3. Proof of Proposition 2
$\log\left(\frac{\pi_{k,g,t}^{0}}{\pi_{p,g,t}^{0}}\right)=\log\left(\frac{q_{k,g,t}^{0}}{S_{g,t}^{0}}\times\frac{S_{g,t}^{0}}{q_{p,g,t}^{0}}\right)=\log(q_{k,g,t}^{0})-\log(q_{p,g,t}^{0}).$
@@
$\log(q_{p,1,1}^{0})-\log(q_{p,1,0}^{0})=\log(q_{p,0,1}^{0})-\log(q_{p,0,0}^{0}).$
Taking the difference between the two equations gives:
$\big\{\log(q_{k,1,1}^{0})-\log(q_{p,1,1}^{0})\big\}-\big\{\log(q_{k,1,0}^{0})-\log(q_{p,1,0}^{0})\big\}=$
$\big\{\log(q_{k,0,1}^{0})-\log(q_{p,0,1}^{0})\big\}-\big\{\log(q_{k,0,0}^{0})-\log(q_{p,0,0}^{0})\big\}$
@@
$\log\left(\frac{\pi_{k,1,1}^{0}}{\pi_{l,1,1}^{0}}\right)-\log\left(\frac{\pi_{k,0,1}^{0}}{\pi_{l,0,1}^{0}}\right)=\log\left(\frac{\pi_{k,1,0}^{0}}{\pi_{l,1,0}^{0}}\right)-\log\left(\frac{\pi_{k,0,0}^{0}}{\pi_{l,0,0}^{0}}\right).$
From part 1, I can also derive that:
$\big\{\log(q_{k,1,1}^{0})-\log(q_{p,1,1}^{0})\big\}-\big\{\log(q_{k,1,0}^{0})-\log(q_{p,1,0}^{0})\big\}=$
$\big\{\log(q_{k,0,1}^{0})-\log(q_{p,0,1}^{0})\big\}-\big\{\log(q_{k,0,0}^{0})-\log(q_{p,0,0}^{0})\big\}$
@@
$\Big\{\mathbb{E}[U^{0}_{k,1,1}]-\mathbb{E}[U^{0}_{11}]\Big\}-\Big\{\mathbb{E}[U^{0}_{k,1,0}]-\mathbb{E}[U^{0}_{k,1,0}]\Big\}=$
$\Big\{\mathbb{E}[U^{0}_{k,0,1}]-\mathbb{E}[U^{0}_{k,0,1}]\Big\}-\Big\{\mathbb{E}[U^{0}_{k,0,0}]-\mathbb{E}[U^{0}_{l,0,0}]\Big\},$ (15)